OpenStudy (anonymous):

ok help me out again... 1/25=125^8x+3

OpenStudy (anonymous):

5^(-2) = 5^(24x+9) -2 = 24x +9 -11 = 24x x = -11/24

OpenStudy (angela210793):

125^8x+3=5^3(8x+3) 1/25=5^-2 then as Josephniel :)